qt=r(s+t). Solve for t.
The answer is supposed to be t=rs over q-r I can't find a way to get the correct answer. Help!

Answer by richwmiller(17219)

qt=r(s+t)
multiply out
qt=rs+rt
subtract rt
qt-rt=rs
factor out t
t*(q-r)=rs
divide by q-r
t=rs/(q-r)